# FSDAM: Few-Shot Driver Attention Modeling via Vision-Language Coupling

### v1

@@ Abstract
Understanding where drivers look and why they shift their attention is essential for autonomous systems that read human intent and justify their actions. Most existing models rely on large-scale gaze datasets to learn these patterns; however, such datasets are labor-intensive to collect and time-consuming to curate. We present FSDAM (Few-Shot Driver Attention Modeling), a framework that achieves joint attention prediction and caption generation with approximately 100 annotated examples, two orders of magnitude fewer than existing approaches. Our approach introduces a dual-pathway architecture where separate modules handle spatial prediction and caption generation while maintaining semantic consistency through cross-modal alignment. Despite minimal supervision, FSDAM achieves competitive performance on attention prediction, generates coherent, and context-aware explanations. The model demonstrates robust zero-shot generalization across multiple driving benchmarks. This work shows that effective attention-conditioned generation is achievable with limited supervision, opening new possibilities for practical deployment of explainable driver attention systems in data-constrained scenarios.
@@ 1 Introduction
Driving is primarily a visual task. Drivers constantly move their gaze across different environmental elements, such as traffic signals, crossing pedestrians, and road hazards. This selective attention directly influences their understanding of the situation and decision-making process. Suboptimal attention allocation due to distraction or drowsiness remains a leading cause of accidents, with disproportionate crash rates at intersections where drivers fail to notice crossing traffic, or during lane changes when blind spots go unchecked [9312486, 10588743]. Understanding where and what drivers visually attend to can also be beneficial to Advanced Driver Assistance Systems (ADAS) and autonomous vehicles (AVs) operating in mixed-traffic environments, where human-AI collaboration depends on a shared perception of the surroundings. Modeling driver attention allows for cross-verification of the human-machine goal alignment by comparing human intention with decisions made by autonomous systems. This process validates whether the planned maneuvers are consistent with the perceived context of the scene, and it also supports interpretable decision-making chains that enhance user trust in human-AI collaboration.
Figure 1: Few-shot learning performance on BDD-A dataset. Gaze prediction metrics: $C$, SIM, NSS, AUC-J, AUC-B, and KL (inverted). All metrics are linearly min-max normalized to [0,1] where larger area indicates better performance.
Despite its importance, effectively modeling driver attention faces fundamental challenges. Existing approaches rely heavily on supervised learning that requires large, labeled eye-tracking data collected in well-represented conditions to provide stable inductive biases for low-level perception. However, the data collection and annotation can be costly and often lack diversity in safety-critical scenarios [li2025recogdrive]. Additionally, an image dataset containing driver gaze information can expose sensitive biometric and behavioral data, potentially revealing individuals’ identity and attention patterns [Kröger2020]. This data scarcity, coupled with privacy and annotation costs, can lead to long-tail errors during domain-specific corner cases and rare events [ghosh2025roadwork]. Moreover, the space of possible driving scenarios grows combinatorially with environmental factors, traffic configurations, and agent behaviors, making comprehensive coverage through dataset expansion increasingly impractical [ghosh2025roadwork, li2025recogdrive]. These constraints necessitate a shift from data-hungry supervised learning to sample-efficient learning with semantic grounding. In addition to learning visual attention patterns through labeled datasets, semantic understanding of the surrounding environment can also benefit visual attention prediction. For example, knowing that “pedestrians near crosswalks" or “vehicles braking suddenly" are situations where driver attention is typically expected, and can be encoded as an auxiliary loss that encourages the model’s attention distribution to place more weight on these regions. Vision-language models (VLMs) can offer this capability: their pre-trained representations encode rich semantic knowledge about objects, spatial relationships, and causal reasoning. Rather than learning attention patterns from scratch, VLMs can exploit these semantic priors to achieve few-shot generalization with as few as 5-32 examples per class [radford2021learning]. In autonomous driving, works like DriveLM [sima2023drivelm] have demonstrated that grounding perception in natural language enables more interpretable reasoning that generalizes across scenarios. Building on this insight, Zhou et al. [zhou2025where] pioneered joint attention-language modeling with LLaDA, demonstrating that coupling spatial attention prediction with textual reasoning improves both quantitative performance and interpretability. Similarly, Chen et al. [chen2024gazexplainlearningpredictnatural] showed that generating natural language explanations enhances gaze prediction understanding. While these works demonstrate the potential of vision-language models for joint spatial and semantic reasoning, they still require extensive labeling. The objective of this work is therefore to leverage VLM priors to learn driver attention from fewer labeled examples.
@@
To address these limitations, we develop FSDAM, a framework that learns joint gaze prediction and attention-grounded caption generation from approximately 100 annotated examples. To our knowledge, this is the first work to demonstrate that attention-conditioned language generation can be learned effectively in a few-shot regime. Beyond data efficiency, we introduce attention shift anticipation: unlike prior work that describes current attention, our model predicts where drivers will look next from single frames.c We create supervision by selecting frame pairs where attention demonstrably shifts and annotating observed transitions with structured captions describing scene context, current focus, anticipated shift, and causal reasoning. At inference, the model learns to anticipate these shifts from single frames alone, capturing the forward-looking nature of human attention without requiring video input. Our key insight is to structure the learning task so that spatial attention and language understanding provide complementary supervision signals. We introduce a structured attention reasoning representation where each annotation describes the driving scene, identifies where the driver is currently looking, predicts where attention will shift next, and explains why. This structured representation creates explicit connections between visual attention patterns and semantic reasoning.
Our architecture processes visual features through two pathways: one generates attention-grounded captions while the other aligns predicted attention with language representations. Both pathways share the same visual features but specialize for their respective tasks. Starting from approximately 100 training examples curated from the Berkeley DeepDrive Attention (BDD-A) dataset [10.1007/978-3-030-20873-8_42], our model achieves strong spatial attention prediction despite the severe data limitation. It reduces KL divergence on BDD-A by up to 89% compared to fully supervised baselines and delivers competitive CC and NSS values. Zero shot transfer experiments on DR(eye)VE [palazzi2018predicting] and DADA 2000 [fang2023dadadriverattentionprediction] further show that FSDAM maintains the lowest KL divergence and strong SIM scores across these unseen domains, indicating that the learned attention language alignment generalizes well beyond its training distribution. This efficient learning regime supports practical use in driving environments where gaze data is limited, such as night scenes, weather affected roads, and high speed highways. In summary, we make the following contributions:
@@
  • First few-shot approach for attention-based generation. We are the first of its kind to achieve joint spatial attention prediction and natural language explanation in a few-shot learning regime, training from fewer than 100 examples, two orders of magnitude more data efficient than existing joint modeling approaches [zhou2025where, chen2024gazexplainlearningpredictnatural].
  • Attention shift anticipation. Instead of focusing on current attention, our model introduces a structured attention reasoning representation that captures scene context, current focus, anticipated shift and causal factors. This enables models to predict where attention will move next from single frames, capturing the forward-looking nature of human attention without requiring video input at inference.
  • Dual-pathway architecture for joint learning. We design an architecture with separate pathways for caption generation and attention-language alignment that share visual representations, enabling task-specific specialization while maintaining consistency.
  • Strong empirical validation. Despite training in few-shot regime, we achieve competitive gaze prediction and caption generation with strong zero-shot transfer across three standard benchmarks demonstrating generalization beyond the training data distribution.
@@ 2.2 Vision-Language Models for Explainable Attention
For driver attention specifically, LLada [zhou2025where] proposes joint modeling of attention maps and textual descriptions through their W3DA dataset ( 70k samples), predicting both where drivers look and why attention is allocated. GazeXplain [chen2024gazexplainlearningpredictnatural] generates natural language descriptions of gaze scanpaths for general visual attention.
@@ 2.3 Few-Shot Learning and Data-Efficient Adaptation
Figure 2: Dataset curation pipeline. From BDD-A [10.1007/978-3-030-20873-8_42] videos to paired frames, GPT-4o captioning with fixed template, human verification, and final captions. Full details in Appendix.
@@ 3 Method
We present our few-samples dependent driver attention modeling framework that joints modeling of spatial driver attention prediction and structured natural language explanation. These two tasks depend on totally different kinds of spatial understanding, where captioning needs global semantic reasoning while gaze prediction requires localized spatial sensitivity. A shared cross-attention module tends to overfit one task and underfit another [Li_2024, fifty2021efficientlyidentifyingtaskgroupings]. To address this imbalance, we introduce a dual-pathway architecture (Figure 3) in which gaze prediction and caption generation are handled by separate modules while still leveraging shared visual features. A vision-language alignment mechanism further provides semantic supervision to spatial prediction, ensuring predicted attention regions correspond to meaningful visual content. This design enables effective joint learning despite data scarcity.
Figure 3: FSDAM architecture for joint gaze prediction and caption generation. A frozen vision-language backbone extracts spatial features $F_{t}$ and text embeddings $q_{\text{text}}$. The gaze pathway, left, upsamples $F_{t}$ to predict attention and gaze maps $\widehat{G}$. The caption pathway, right, uses cross-attention over $F_{t}$ to generate structured explanations.
@@ 3.1 Problem Formulation
Here, $C_{{scene}}$ summarizes the global scene context, $C_{current}$ describes the driver’s present focus of attention, $C_{next}$ anticipates how the attention is likely to shift next (“will check the crosswalk”), and $C_{why}$ provides the causal rationale underlying this transition. Together, the four components form a structured attention reasoning representation that links spatial attention patterns with high-level semantic explanations, enabling the model to learn gaze behavior from single frames.
Our architecture builds upon LLaVA-Next-1.6 [liu2024llavanext], which comprises a CLIP-ViT-L/14 vision encoder [radford2021learning] extracting spatial features $F_{t}\in\mathbb{R}^{B\times 1024\times 24\times 24}$ and a Mistral-7B language model [jiang2023mistral7b] producing image-conditioned text embeddings $q_{\text{text}}\in\mathbb{R}^{B\times d_{\ell}}$ through multimodal fusion [alayrac2022flamingo, liu2024llavanext]. These frozen features feed into specialized pathways for gaze prediction and caption generation, with task-specific adaptation enabled by LoRA [hu2022lora]. We optimize 13.6M task-specific parameters (0.2% of the 7B backbone) while keeping pretrained components frozen to prevent overfitting.
@@ 3.2 Spatial Attention Prediction
The gaze pathway predicts where drivers look. A convolutional decoder upsamples $F_{t}$ from $24{\times}24$ to $64{\times}64$ resolution through successive convolution and bilinear upsampling layers [long2015fullyconvolutionalnetworkssemantic, ronneberger2015unetconvolutionalnetworksbiomedical], then applies spatial softmax to produce $\widehat{G}\in\Delta^{S\times S}$.
We supervise with forward KL divergence to encourage covering all human fixation regions:
@@
This alone produces overly diffuse predictions under limited data. We augment with a blur-gap regularizer that applies Gaussian smoothing ($\sigma{=}1.0$) to $\widehat{G}$ to obtain $\widetilde{G}$, then penalizes cases where blurring doesn’t degrade prediction:
@@
with $\lambda{=}0.3$ and $\epsilon{=}0.05$. This encourages confident spatial localization in low-data regimes [kümmerer2016deepgazeiireadingfixations, cornia2018predicting]. The gaze pathway receives gradients only from $\mathcal{L}_{\text{gaze}}$ during forward passes, with semantic supervision added through alignment (Section 3.4).
@@ 3.3 Attention-Grounded Caption Generation
where $y$ contains scene description, current attention, anticipated shift, and causal reasoning. LoRA [hu2022lora] enables task-specific tuning while keeping most parameters frozen.
@@ 3.4 Vision-Language Alignment
using the model’s own prediction maintains consistency between training and inference [chen2020simpleframeworkcontrastivelearning, grill2020bootstrap].
@@ 3.5 Training and Inference
Training.
We train using AdamW optimizer with separate learning rates for LoRA and task modules. Training uses mixed precision, gradient accumulation for an effective batch of 16, and input size $336{\times}336$. Only LoRA and lightweight adapters are updated, totaling 13.6M trainable parameters ($\sim$0.2% of the backbone). Learning rate is adjusted via a ReduceLROnPlateau scheduler.
Inference.
During inference, only gaze and caption branches are active. The gaze decoder predicts $\widehat{G}$ from visual features $F_{t}$, while the caption pathway generates structured explanations through cross-attention and prefix adaptation. The alignment module is used only during training and adds no cost at inference.
@@ 4 Experimental Setup
Dataset Preparation.
We construct a gaze-language dataset from BDD-A dataset [10.1007/978-3-030-20873-8_42], which contains braking event videos selected from large-scale, crowd-sourced driving video data combined with compiled human gaze data from 6.5 seconds prior and 3.5 seconds after each braking event. To identify frames capturing meaningful attention transitions, we employ a KL-divergence-based selection algorithm [shlens2014noteskullbackleiblerdivergencelikelihood] that detects moments of maximum gaze distribution change between consecutive frames. Local peaks in the KL divergence curve mark anchor frames $t$ where attention shifts abruptly. For each anchor, we select a target frame $t+\Delta$ within $[3,18]$ frames that maximizes divergence from the anchor, capturing the strongest future attention transition. Following temporal sampling strategies from video action recognition [carreira2018quovadisactionrecognition], we filter clips shorter than 50 frames and retain the top - K = 2 anchor–target pairs per video for temporal diversity. Each resulting sample $(I_{t},I_{t+\Delta},G_{t},G_{t+\Delta})$ consists of two frames and their corresponding gaze maps. We generate structured captions via GPT-4o [openai2024gpt4technicalreport] with human verification (Figure 2), producing 90 training examples across eight diverse driving scenarios.
Implementation Details.
All experiments are implemented in PyTorch [paszke2019pytorchimperativestylehighperformance]. We build on the LLaVA-Next Mistral-7B backbone, which integrates a CLIP-based vision tower pretrained on large-scale image–text pairs. The vision–language backbone remains frozen during training, and only lightweight task-specific components (gaze decoder, prefix adapter, alignment head, and LoRA adapters) are updated. Training is performed on a single NVIDIA GH200 GPU with input resolution $336{\times}336$, batch size 4, and gradient accumulation over 4 steps (effective batch 16). We use the AdamW optimizer with learning rates of $1{\times}10^{-4}$ for LoRA and $2{\times}10^{-4}$ for task heads, a ReduceLROnPlateau scheduler (factor 0.5, patience 2), and mixed precision of bfloat16.
Evaluation Metrics
We evaluate our model using standard saliency and captioning metrics. For gaze prediction, we adopt CC, KL, SIM, AUC-J, AUC-B, and NSS following the official MIT Saliency Benchmark [bylinskii2017differentevaluationmetricstell]. For caption generation, we report BLEU, METEOR, ROUGE-L, CIDEr-R, and BERTScore using the COCO Caption Evaluation Toolkit [chen2015microsoftcococaptionsdata]. All metrics are computed using their standardized implementations to ensure fair comparison with prior work.

### v2

@@ Abstract
Understanding not only where drivers look but also why their attention shifts is essential for interpretable human–AI collaboration in autonomous driving. Driver attention is not purely perceptual but semantically structured. Thus attention shifts can be learned through minimal semantic supervision rather than dense large-scale annotation. We present FSDAM (Few-Shot Driver Attention Modeling), a framework that achieves joint spatial attention prediction and structured explanation generation using 90 annotated examples. Our key insight is to decompose attention into an explicit reasoning representation, including scene context, current focus, anticipated next focus, and causal explanation, and to learn next-focus anticipation through minimal-pair supervision. To address task conflict and large sample requirements of existing models, and to mitigate task interference under limited data, we introduce a novel dual-pathway architecture in which separate modules handle spatial prediction and caption generation. In addition, we use a training-only vision–language alignment mechanism that injects semantic priors into spatial learning without increasing inference complexity, mitigating task interference under few-shot training. Despite extreme data scarcity, FSDAM achieves competitive performance in gaze prediction, and generates coherent, context-aware structural reasoning for improved interpretability. The model further demonstrates strong zero-shot generalization across multiple driving benchmarks. These results suggest that semantically grounded attention modeling enables data-efficient learning and provides a scalable path toward explainable driver attention systems in data-constrained environments.
@@ 1 Introduction
Figure 1: Few-Shot Driving Attention Modeling (FSDAM). (A) Joint prediction of a gaze map and a structured reasoning (Scene, Current Gaze, Next Gaze, Why). (B) Key frame selection on BDD-A via KL-divergence mining to select high-change gaze-transition moments between frame pairs. (C) Dual-pathway design with separate gaze and caption pathways, coupled through a training-only vision–language alignment loss. (D) Qualitative gaze predictions across four representative driving scenarios.
Driving is a fundamentally visual and anticipatory task. Drivers continuously allocate attention across traffic signals, pedestrians, vehicles, and road geometry, not only passively reacting to static observation but actively anticipating potential hazards and right-of-way conflicts. This selective attention critically shapes situation awareness (SA) and downstream decision-making. Suboptimal attention allocation with critical actors overlooked remains a major contributor to traffic accidents, with high crash rates at intersections and during lane changes [9312486, 10588743]. Understanding where and why drivers allocate attention is therefore essential not only for safety analysis but also for autonomous vehicles (AVs) operating in mixed-traffic environments. Furthermore, modeling driver attention enables cross-verification between human intent and machine decisions in human-AI collaboration, supporting interpretable and trustworthy autonomy.
Despite its importance, modeling driver attention remains challenging. Most existing approaches formulate attention prediction as a supervised saliency estimation problem, learning spatial gaze distributions from large-scale eye-tracking datasets. These methods depend heavily on densely annotated data collected under well-represented conditions to stabilize low-level perception [li2025recogdrive]. However, large-scale gaze data collection is costly, privacy-sensitive, and often lacks diversity in safety-critical or rare scenarios [kroger2020does, ghosh2025roadwork]. Moreover, the combinatorial growth of traffic configurations makes exhaustive dataset expansion increasingly impractical [ghosh2025roadwork, li2025recogdrive]. This reliance on extensive supervision limits scalability and generalization to new domains.
Beyond data scarcity, a more fundamental limitation lies in how attention is conceptualized. Treating driver attention as static perceptual saliency overlooks its inherently forward-looking and semantically structured nature. Human drivers do not simply fixate on visually prominent regions; they anticipate future events based on traffic rules, agent behaviors, and scene context. Recent vision–language models (VLMs) offer a promising direction by injecting semantic reasoning into attention modeling. Works such as DriveLM [sima2023drivelm], LLada [zhou2025where], and GazeXplain [chen2024gazexplainlearningpredictnatural] demonstrate that coupling spatial attention with language reasoning improves interpretability and performance. However, these approaches typically require tens of thousands of annotated samples to jointly align gaze and language representations, limiting their applicability in data-constrained scenarios.
@@
To address these limitations, we develop FSDAM (Figure 1), a framework that jointly learns gaze prediction and attention-grounded caption generation from 90 annotated examples. To our knowledge, this is the first work showing that attention-conditioned language generation can be learned effectively in a few-shot regime. We also introduce a next-attention target prior derived from semantic cues via minimal-pair supervision. Unlike motion-based temporal forecasting, our method estimates future gaze positions from a single frame, conditioned on the current scene. We construct supervision by selecting frame pairs with clear attention shifts and annotating each transition with a structured caption describing scene context, current focus, anticipated next focus, and causal rationale. At inference, the model captures the forward-looking nature of human attention without requiring video input. Our key insight is to couple spatial attention and language understanding as complementary supervision signals through an explicit, structured attention-reasoning format. Building on this design, we propose a dual-pathway architecture that decouples gaze prediction from caption generation, enabling effective joint learning under extreme data scarcity. In summary, we make the following contributions:
@@
  • First few-shot approach for attention-based generation. We are the first of its kind to achieve joint spatial attention prediction and natural language explanation in a few-shot learning regime, training from 90 examples, two orders of magnitude more data efficient than existing joint modeling approaches [zhou2025where, chen2024gazexplainlearningpredictnatural].
  • Structured attention reasoning formulation. We reformulate driver attention modeling as structured anticipation rather than static saliency prediction. Specifically, we decompose attention into four explicit components, including scene context, current focus, anticipated next focus, and causal explanation, establishing a semantic representation that links spatial gaze patterns with forward-looking reasoning.
  • Dual-pathway architecture with training-only alignment. We propose a decoupled gaze–language architecture that mitigates negative transfer under limited data. A training-only vision–language alignment mechanism injects semantic priors into spatial prediction without increasing inference complexity.
  • Data-efficient and transferable attention modeling. Despite training in a few-shot regime, our method achieves competitive performance against fully supervised baselines trained on more data and demonstrates strong zero-shot generalization across multiple driver attention benchmarks.
@@ 2.2 Vision-Language Models for Explainable Attention
For driver attention specifically, LLada [zhou2025where] proposes joint modeling of attention maps and textual descriptions through their W3DA dataset ($\sim$70k samples), predicting both where drivers look and why attention is allocated. GazeXplain [chen2024gazexplainlearningpredictnatural] generates natural language descriptions of gaze scanpaths for general visual attention.
@@ 2.3 Few-Shot Learning and Data-Efficient Adaptation
Figure 2: Dataset curation pipeline. From BDD-A [10.1007/978-3-030-20873-8_42] videos to paired frames, GPT-4o captioning with fixed template, human verification, and final captions. Full details in supplementary material.
@@ 3 Method
We present our few-shot dependent driver attention modeling framework that jointly models spatial driver attention prediction and structured natural language explanation. These two tasks depend on totally different kinds of spatial understanding, where captioning needs global semantic reasoning while gaze prediction requires localized spatial sensitivity. A shared cross-attention module tends to overfit one task and underfit another [Li_2024, fifty2021efficientlyidentifyingtaskgroupings]. To address this imbalance, we introduce a dual-pathway architecture (Figure 3) in which gaze prediction and caption generation are handled by separate modules while still leveraging shared visual features. A vision-language alignment mechanism further provides semantic supervision to spatial prediction, ensuring predicted attention regions correspond to meaningful visual content. This design enables effective joint learning despite data scarcity.
Figure 3: Overview of the proposed FSDAM architecture (right). A frozen vision–language backbone extracts spatial features $F_{t}$ and text embeddings $q_{\text{text}}$. The gaze pathway predicts the attention map $\widehat{G}$, while the explanation pathway performs cross-attention over $F_{t}$ to generate structured reasoning. The training-only vision–language alignment module (top left) applies contrastive supervision between gaze-weighted pooled visual features and text features, and backpropagates gradients to the gaze decoder. The gaze decoder head (bottom left) upsamples $F_{t}$ to produce $\widehat{G}$.
@@ 3.1 Problem Formulation
Here, each $C_{*}$ is a natural language sentence describing one aspect of the attention state. Specifically, $C_{{scene}}$ summarizes the global scene context, $C_{current}$ describes the driver’s present focus of attention, $C_{\text{next}}$ specifies the likely next focus of attention (“will check the crosswalk”), and $C_{why}$ provides the causal rationale underlying this transition. Together, the four components form a structured attention reasoning representation that links spatial attention patterns with high- level semantic explanations, enabling the model to learn gaze behavior from single frames.
Our architecture builds upon LLaVA-Next-1.6 [liu2024llavanext], which comprises a CLIP-ViT-L/14 vision encoder [radford2021learning] extracting spatial features $F_{t}\in\mathbb{R}^{B\times 1024\times 24\times 24}$ for input frame $t$, and a Mistral-7B language model [jiang2023mistral7b] producing image-conditioned text embeddings $q_{\text{text}}\in\mathbb{R}^{B\times d_{\ell}}$, where $d_{\ell}$ is the language model’s hidden dimension, through multimodal fusion [alayrac2022flamingo, liu2024llavanext]. These frozen features feed into specialized pathways for gaze prediction and caption generation, with task-specific adaptation enabled by LoRA [hu2022lora]. We optimize 13.6M task-specific parameters (0.2% of the 7B backbone) while keeping pretrained components frozen to prevent overfitting.
@@ 3.2 Spatial Attention Prediction
The gaze pathway predicts where drivers look. A convolutional decoder upsamples $F_{t}$ from $24{\times}24$ to $64{\times}64$ resolution through successive convolution and bilinear upsampling layers [long2015fullyconvolutionalnetworkssemantic, ronneberger2015unetconvolutionalnetworksbiomedical], then applies spatial softmax to produce $\widehat{G}\in\Delta^{S\times S}$. We supervise with forward KL divergence to encourage covering all human fixation regions:
@@
Here, $G$ is the ground-truth gaze distribution and $i,j$ index spatial positions over the $S{\times}S$ grid. This alone produces overly diffuse predictions under limited data. We augment with a blur-gap regularizer that enforces spatial sharpness by penalizing predictions whose quality does not degrade under Gaussian smoothing ($\sigma{=}1.0$). Specifically, we smooth $\widehat{G}$ to obtain $\widetilde{G}$, then penalize cases where blurring does not degrade the prediction:
@@
with $\lambda{=}0.3$ and $\epsilon{=}0.05$. This encourages confident spatial localization by exploiting the sharpness–blur relationship in saliency maps [kümmerer2016deepgazeiireadingfixations], where diffuse predictions degrade minimally under Gaussian smoothing while sharp, well-localized predictions do not. The gaze pathway receives gradients only from $\mathcal{L}_{\text{gaze}}$ during forward passes, with semantic supervision added through alignment (Section 3.4).
@@ 3.3 Attention-Grounded Caption Generation
where $y$ contains scene description, current attention, next focus of attention, and causal reasoning. LoRA [hu2022lora] enables task-specific tuning while keeping most parameters frozen.
@@ 3.4 Vision-Language Alignment
Here, using the model’s own predicted attention as pooling weights maintains consistency between training and inference [chen2020simpleframeworkcontrastivelearning, grill2020bootstrap].
@@ 3.5 Training and Inference
Training. We train using AdamW optimizer with separate learning rates for LoRA and task modules. Training uses mixed precision, gradient accumulation for an effective batch of 16, and input size $336{\times}336$. Only LoRA and lightweight adapters are updated, totaling 13.6M trainable parameters ($\sim$0.2% of the backbone). Learning rate is adjusted via a ReduceLROnPlateau scheduler.
Inference. During inference, only gaze and caption branches are active. The gaze decoder predicts $\widehat{G}$ from visual features $F_{t}$, while the caption pathway generates structured explanations through cross-attention and prefix adaptation. The alignment module is used only during training and adds no cost at inference.
@@ 4 Experimental Setup
Dataset Preparation. We construct a gaze-language dataset from BDD-A dataset [10.1007/978-3-030-20873-8_42], which contains braking event videos selected from large-scale, crowd-sourced driving video data combined with compiled human gaze data from 6.5 seconds prior and 3.5 seconds after each braking event. To identify frames capturing meaningful attention transitions, we employ a KL-divergence-based selection algorithm [shlens2014noteskullbackleiblerdivergencelikelihood] that detects moments of maximum gaze distribution change between consecutive frames. Local peaks in the KL divergence curve mark anchor frames $t$ where attention shifts abruptly. For each anchor, we select a target frame $t+\Delta$ within $[3,18]$ frames that maximizes divergence from the anchor, capturing the strongest future attention transition. Following temporal sampling strategies from video action recognition [carreira2018quovadisactionrecognition], we filter clips shorter than 50 frames and retain the top-K = 2 anchor– target pairs per video for temporal diversity. Each resulting sample $(I_{t},I_{t+\Delta},G_{t},G_{t+\Delta})$ consists of two frames and their corresponding gaze maps. We generate structured captions via GPT-4o [openai2024gpt4technicalreport] with human verification (Figure 2), yielding 90 training examples systematically sampled from the candidate key frame pairs to evenly cover eight driving scenario categories (see supplementary for the taxonomy and category-wise counts).
Implementation Details. All experiments are implemented in PyTorch [paszke2019pytorchimperativestylehighperformance]. We build on the LLaVA-Next Mistral- 7B backbone, which integrates a CLIP-based vision tower pretrained on large-scale image–text pairs. The vision–language backbone remains frozen during training, and only lightweight task-specific components (gaze decoder, prefix adapter, alignment head, and LoRA adapters) are updated. Training is performed on a single NVIDIA GH200 GPU with input resolution $336{\times}336$, batch size 4, and gradient accumulation over 4 steps (effective batch 16). We use the AdamW optimizer with learning rates of $1{\times}10^{-4}$ for LoRA and $2{\times}10^{-4}$ for task heads, a ReduceLROnPlateau scheduler (factor 0.5, patience 2), and mixed precision of bfloat16.
Evaluation Metrics. We evaluate our model using standard saliency and captioning metrics. For gaze prediction, we adopt CC, KL, SIM, AUC-J, AUC-B, and NSS following the official MIT Saliency Benchmark [bylinskii2017differentevaluationmetricstell]. For caption generation, we report BLEU, METEOR, ROUGE-L, CIDEr-R, and BERTScore using the COCO Caption Evaluation Toolkit [chen2015microsoftcococaptionsdata]. All metrics are computed using their standardized implementations to ensure fair comparison with prior work.